Solve the following system by substitution.
9x-y=10
y=-9x+6
Could you send the solution? Thanks.

Answer by stanbon(75887)

Solve the following system by substitution.
9x-y=10
y=-9x+6
-----
Substitute for "y"; solve for "x"::
9x - (-9x+6) = 10
------
9x + 9x - 6 = 10
-----
18x = 16
x = 8/9
-----------------
Solve for "y"::
y = -9(8/9)+6
y = -8 + 6
y = -2
------------
Cheers,
Stan H.
------------